OpenStudy (anonymous):

Can you justify if completing the square is a good method for solving when the Discriminant is negative? Use any of these three functions as an example f(x)=x^2-4 g(x)=x^2-2 h(x)=-3x^2 +11i

OpenStudy (amistre64):

completing the square is the proof of the quadratic formula.

OpenStudy (amistre64):

\[ax^2+bx+c=0\] \[a(x^2+\frac{b}{a}x)+c=0\] \[a(x^2+\frac{b}{a}x\pm \frac{b^2}{4a^2})+c=0\] \[a(x^2+\frac{b}{a}x+\frac{b^2}{4a^2})-\frac{b^2}{4a}+c=0\] \[a(x+\frac{b}{2a})^2-\frac{b^2}{4a}+c=0\] \[a(x+\frac{b}{2a})^2=\frac{b^2}{4a}-c\] \[(x+\frac{b}{2a})^2=\frac{b^2}{4a^2}-\frac ca\] \[(x+\frac{b}{2a})^2=\frac{b^2}{4a^2}-\frac {4ac}{4aa}\] \[x+\frac{b}{2a}=\pm\sqrt{\frac{b^2}{4a^2}-\frac {4ac}{4a^2}}\] \[x+\frac{b}{2a}=\pm\frac{\sqrt{b^2-4ac}}{2a}\] \[x=-\frac{b}{2a}\pm\frac{\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

what?

OpenStudy (amistre64):

like i said: completing the square is the proof of the quadratic formula. since the quadratic formula provides solutions; then so does completing the square ... they are 2 sides of the same coin :/

OpenStudy (anonymous):

do you know how to convert f(x) into general vertex form?